# Unified description of viscous, viscoelastic, or elastic thin active films on substrates

### v1

@@ V Rescaled dynamic equations
$\partial_{\tilde{t}}\mathbf{P}+\tilde{\mathbf{v}}\cdot\tilde{\nabla}\mathbf{P}=$$\,\tilde{\nabla}^{2}\mathbf{P}-\mathbf{P}+\tilde{\boldsymbol{\Omega}}\cdot\mathbf{P}+\kappa\tilde{\boldsymbol{\Sigma}}\cdot\mathbf{P}$ (23)
@@
In terms of the dimensionless displacement field $\tilde{\mathbf{u}}=\mathbf{u}/\sqrt{D_{\mathrm{tr}}/D_{\mathrm{r}}}$, the evolution equation for the velocity field, Eq. (18), becomes
@@
The rescaled pressure $\tilde{p}=D_{\mathrm{r}}^{-1}D_{\mathrm{tr}}^{-1}\rho^{-1}p$ acts as a Lagrange multiplier ensuring incompressibility via a divergence-free velocity field, ~$\tilde{\nabla}\cdot\tilde{\mathbf{v}}=0$. Finally, we recast Eq. (20) for the evolution of the displacement field into the dimensionless form
@@ VI Analysis for spatially uniform solutions
We begin with stationary solutions. Thus, we set all time derivatives to zero, which yields
@@
always exists, regardless of the chosen parameter values. It describes an isotropic state of disordered swimmer orientations, vanishing macroscopic velocity (macroscopic quiescence), and vanishing deformation.
@@
$v_{0}^{\mathrm{st}}$ $=\frac{\nu_{\mathrm{p}}}{\nu_{\mathrm{v}}+\nu_{\mathrm{d}}\tau_{\mathrm{d}}}P_{0}\,,$
$u_{0}^{\mathrm{st}}$ $=\frac{\nu_{\mathrm{p}}\tau_{\mathrm{d}}}{\nu_{\mathrm{v}}+\nu_{\mathrm{d}}\tau_{\mathrm{d}}}P_{0}\,.$
@@
Figure 2: Spatially uniform dynamic states of rotating polar orientational order and flow. (a) Snapshots of the uniform velocity field during one period $T$ for a genuinely elastic solid ($\tau_{\mathrm{d}}\rightarrow\infty$). The arrows in the main plots display the current state of the overall flow field $\mathbf{v}$. Circular insets visualize the current states of the polar orientational order parameter $\mathbf{P}$, overall flow velocity $\mathbf{v}$, and displacement field $\mathbf{u}$. Phase shifts between these quantities are obvious. These results agree with numerical solutions in a small periodic box. (b) Corresponding time evolution of the spatially averaged $x$-components of $\mathbf{P}$, $\mathbf{v}$, and $\mathbf{u}$. We rescale $\langle v_{x}\rangle$ and $\langle u_{x}\rangle$ by the amplitudes $v_{0}$ and $u_{0}$, respectively. We set the remaining parameters to $\gamma_{\mathrm{a}}=1$, $\kappa=0$, $\eta=1$, $\mu=1$, $\nu_{\mathrm{v}}=1$, $\nu_{\mathrm{d}}=10$, and $\nu_{\mathrm{p}}=20$.
@@
In addition to these stationary solutions, we find analytical, spatially homogeneous solutions of rotating polar order and flow over time. Here, we return to Eqs. (23)–(25) and only set all spatial derivatives to zero. Since Eq. (24) is linear and does not explicitly depend on time, the velocity $\mathbf{v}$ is not an independent quantity. Furthermore, the incompressibility conditions are fulfilled automatically. The pressure field does not enter the solutions, because we are only considering spatially homogeneous situations. The evolution equations thus reduce to
@@
We write polar orientational order parameter, velocity, and displacement fields in polar components,
@@
To summarize, these solutions represent rotating spatially uniform translational polar orientational order, overall flows, and displacements. The magnitudes of polar orientational order, velocity, and displacement fields are constant. Yet, their directions rotate with locked phase shifts, following each other. The sense of rotation results from spontaneous symmetry breaking.
@@ VII Transitions between the different analytical solutions
Mainly, we focus on the nature of the transitions between the different dynamic states. In fact, the rotating polar solution finally emerges when increasing the strength $\nu_{\mathrm{p}}$ of active forcing from the trivial isotropic solution of vanishing polar orientational order, velocity, and displacement field. To shed some light on this transition, we first consider the example of solid-like materials for $\tau_{\mathrm{d}}\to\infty$. The remaining parameters are set to $\gamma_{\mathrm{a}}=1$, $\kappa=0$, $\eta=1$, $\mu=1$, $\nu_{\mathrm{v}}=1$, and $\nu_{\mathrm{d}}=10$. We find that the rotational polar solution emerges from the isotropic state in a subcritical bifurcation when varying $\nu_{\mathrm{p}}$, see Fig. 3(a). There, we plot the magnitude $v_{0}$ of the velocity as a function of $\nu_{\mathrm{p}}$.
In fact, it turns out that the two solutions in Eqs. (36)–(39) and (43), distinguished by the sign ±$\pm A$, correspond to the stable and unstable analytical solutions associated with this subcritical bifurcation. In Fig. 3(a), stable and unstable branches are marked by solid and dashed lines, respectively. The presence of three branches of solution, one corresponding to the isotropic solution, one to the stable rotating polar solution, and one to the unstable rotating polar solution, indicates hysteretic behavior.
@@
We further investigate the subcritical transition to this rotational state by numerical calculations. To this end, we solve Eqs. (23)–(25) for a small spatially periodic systems, see Appendix C for details. The results are displayed in Fig. 3(a) as the red and blue data points. In the first set of calculations, we start from an isotropic, macroscopically quiescent state. Increasing the active forcing $\nu_{\mathrm{p}}$ (red dots), the system at a certain threshold jumps to the rotational solution of finite amplitudes $P_{0}$, $v_{0}$, and $u_{0}$. However, when starting from an already rotating state and decreasing $\nu_{\mathrm{p}}$ (blue dots), the rotations persist down to a lower value of $\nu_{\mathrm{p}}$. The rotational solution vanishes at finite but smaller amplitudes $P_{0}$, $v_{0}$, and $u_{0}$. This confirms the analytically identified region of hysteresis in a subcritical bifurcation scenario. Either a macroscopically quiescent, isotropic state or a solution of continuously rotating uniform spatial translation develops in this region, depending on the initial conditions.
Next, we generalize our results from purely solid-like to viscoelastic behavior that allows for terminal flow. That is, finite values of the relaxation time $\tau_{\mathrm{d}}$ are considered. For the spatially uniform rotational solution, we find that the rotation frequency $\omega_{0}$ of the polar orientational order parameter, the velocity, and the displacement field depend on the strength $\nu_{\mathrm{p}}$ of active forcing and the relaxation time $\tau_{\mathrm{d}}$ of the memorized displacements. Examples for analytical solutions according to Eqs. (36)–(39) are depicted in Fig. 3(b). We recall that decreasing $\tau_{\mathrm{d}}$ moves the system from rather elastic solid-like towards more fluid-like behavior.
@@
Physically, sufficiently long relaxation times $\tau_{\mathrm{d}}$, that is, more pronounced elasticity, hinder the flow $\mathbf{v}$. The elastic restoring forces are too large in this case to allow for persistent migration in the same direction. As a compromise, rotations occur. Reducing the role of elasticity through decreasing $\tau_{\mathrm{d}}$ facilitates unbounded flows $\mathbf{v}$. Then, instead of rotations, persistent stationary flows combined with polar orientational order are observed.
@@ VIII Conclusions
In the more fluid-like case, we find a spatially uniform state of global polar orientational order, similar to what is observed in conventional theories for nonelastic systems [16, 17, 18]. In contrast to that, solid-like properties prevent persistent global motion in one direction due to the associated elastic restoring forces. There, we observe a spatially uniform state of rotating global polar orientational order. It represents globally uniform motion in one direction, while this direction rotates. When increasing the activity starting from the disordered, isotropic state, these rotations emerge in a subcritical transition exhibiting hysteresis. Our unified description reveals that this hysteretic regime vanishes as the bifurcation changes from sub- to supercritical when decreasing the relaxation time $\tau_{\mathrm{d}}$, that is, turning to more fluid-like behavior in the viscoelastic regime. Remarkably, with increasing activity, the dynamic state can transition from isotropic to stationary polar and then to rotating polar order.
Our description bridges the gap between active fluids and active solids, which have by now mostly been considered in separate frameworks. As pointed out in Ref. 58, there is a critical need for this kind of integrative theories particularly for biological media, as they may change their properties dynamically and continuously. For example, during biofilm formation, bacteria excrete extracellular polymers [5, 39, 38]. Accordingly, the film starts from still fluid-like, but over time changes its rheological properties from viscous and viscoelastic to increasingly elastic. In a settled biofilm, the polymer network forms an elastic matrix and the resulting system may then be described as an active solid [55]. On a continuum-theoretical level, our description covers such transitions between fluid- and solid-like behavior. It shall be applied in future studies accordingly.
@@ Acknowledgements.
Appendix A Elastic stress and displacement field
The relaxation equation for the displacement field, Eq. (20), follows from generalized hydrodynamics [59, 60]. We start from the elastic energy density $\epsilon$ in the Eulerian frame of reference. In terms of the components of the strain tensor $\mathbf{\tilde{U}}$, it reads
$\epsilon=\frac{1}{2}K_{ijkm}\tilde{U}_{ij}\tilde{U}_{km}\,,$ (44)
while the components of the strain tensor in terms of the displacement field $\mathbf{u}$ in the Eulerian frame are given by
$\tilde{U}_{ij}=\frac{1}{2}\big{[}\partial_{x_{i}}u_{j}+\partial_{x_{j}}u_{i}-(\partial_{x_{i}}u_{k})(\partial_{x_{j}}u_{k})\big{]}\,.$ (45)
The fourth-rank tensor $\mathbf{K}$ is the stiffness tensor, which for an isotropic material is of the form

### v2

@@ V Rescaled dynamic equations
$\partial_{\tilde{t}}\mathbf{P}+\tilde{\mathbf{v}}\cdot\tilde{\nabla}\mathbf{P}=$$\,\tilde{\nabla}^{2}\mathbf{P}-\mathbf{P}+\tilde{\bm{\Omega}}\cdot\mathbf{P}+\kappa\tilde{\bm{\Sigma}}\cdot\mathbf{P}$ (22)
@@
In terms of the dimensionless elastic memory field $\tilde{\mathbf{u}}=\mathbf{u}/\sqrt{D_{\mathrm{tr}}/D_{\mathrm{r}}}$, the evolution equation for the velocity field, Eq. (18), becomes
@@
The rescaled pressure $\tilde{p}=D_{\mathrm{r}}^{-1}D_{\mathrm{tr}}^{-1}\rho^{-1}p$ acts as a Lagrange multiplier ensuring incompressibility via a divergence-free velocity field, ~$\tilde{\nabla}\cdot\tilde{\mathbf{v}}=0$. Finally, we recast Eq. (20) for the evolution of the elastic memory field into the dimensionless form
@@ VI Analysis for spatially uniform solutions
We begin with stationary solutions. Thus, we set all time derivatives to zero as well, which yields
@@
always exists, regardless of the chosen parameter values. It describes an isotropic state of disordered orientations of the active units, vanishing macroscopic velocity (macroscopic quiescence), and vanishing deformation.
@@
$v_{0}^{\mathrm{st}}$ $=\frac{\nu_{\mathrm{p}}}{\nu_{\mathrm{v}}+\nu_{\mathrm{d}}\tau_{\mathrm{d}}}P_{0}^{\mathrm{st}}\,,$
$u_{0}^{\mathrm{st}}$ $=\frac{\nu_{\mathrm{p}}\tau_{\mathrm{d}}}{\nu_{\mathrm{v}}+\nu_{\mathrm{d}}\tau_{\mathrm{d}}}P_{0}^{\mathrm{st}}\,.$
@@
Figure 2: Spatially uniform dynamic states of rotating polar orientational order and flow. (a) Snapshots of the uniform velocity field during one period $T$ for a genuinely elastic solid ($\tau_{\mathrm{d}}\rightarrow\infty$). The arrows in the main plots display the current state of the overall flow field $\mathbf{v}$. Circular insets visualize the current states of the polar orientational order parameter $\mathbf{P}$, overall flow velocity $\mathbf{v}$, and field of memorized elastic displacements $\mathbf{u}$. Phase shifts between these quantities are obvious. These results agree with numerical solutions in a small periodic box. (b) Corresponding time evolution of the spatially averaged $x$-components of $\mathbf{P}$, $\mathbf{v}$, and $\mathbf{u}$. We rescale $\langle v_{x}\rangle$ and $\langle u_{x}\rangle$ by the amplitudes $v_{0}$ and $u_{0}$, respectively. We set the remaining parameters to $\gamma_{\mathrm{a}}=1$, $\kappa=0$, $\eta=1$, $\mu=1$, $\nu_{\mathrm{v}}=1$, $\nu_{\mathrm{d}}=10$, and $\nu_{\mathrm{p}}=20$.
@@
To illustrate the condition in Eq. (30), we may introduce a Deborah number $\mathrm{De}$ in the present context. This dimensionless number quantifies the ratio between the elastic relaxation time $\tau_{\mathrm{d}}$ of the viscoelastic medium and some intrinsic time scale $\tau_{a}$ of the considered process, $\mathrm{De}=\tau_{\mathrm{d}}/\tau_{a}$. In our case, the latter is given by the time scale $\tau_{a}$ of active propulsion, and we define it as $\tau_{a}=(2\nu_{\mathrm{p}}\gamma_{\mathrm{a}}/3-\nu_{\mathrm{v}})/\nu_{\mathrm{d}}$. It essentially characterizes the time scale resulting from actively forced, orientationally ordered motion as reduced by linear friction with the substrate when competing with elastic pinning to the substrate. Using these definitions, Eq. (30) reduces to $\mathrm{De}<1$. Smaller Deborah numbers correspond to a situation in which the viscoelastic medium behaves more fluid-like in the context of the considered process. In this case, the process of active propulsion is slower than the relaxation of the elastic memory, which enables the emergence of the stationary polar state and persistent migration. Conversely, for Deborah numbers $\mathrm{De}>1$, the elastic forces prevent persistent directed motion. Thus, the medium behaves more solid-like.
In addition to the two stationary solutions mentioned above, we find analytical, spatially homogeneous solutions of rotating polar order and flow over time. Here, we return to Eqs. (22)–(24) and only set all spatial derivatives to zero. Since Eq. (23) is linear and does not explicitly depend on time, the velocity $\mathbf{v}$ is not an independent quantity. Furthermore, the incompressibility conditions are fulfilled automatically. The pressure field does not enter the solutions, because we are only considering spatially homogeneous situations. The evolution equations thus reduce to
@@
We write polar orientational order parameter, velocity, and elastic memory fields in polar components,
@@
To summarize, these solutions represent rotating spatially uniform translational polar orientational order, overall flows, and displacements. The magnitudes of polar orientational order, velocity, and elastic memory fields are constant. Yet, their directions rotate with locked phase shifts, following each other. The sense of rotation results from spontaneous symmetry breaking.
@@ VII Transitions between the different analytical solutions
Mainly, we focus on the nature of the transitions between the different dynamic states. In fact, the rotating polar solution finally emerges when increasing the strength $\nu_{\mathrm{p}}$ of active forcing from the trivial isotropic solution of vanishing polar orientational order, velocity, and elastic memory field. To shed some light on this transition, we first consider the example of solid-like materials for $\tau_{\mathrm{d}}\to\infty$. The remaining parameters are set to $\gamma_{\mathrm{a}}=1$, $\kappa=0$, $\eta=1$, $\mu=1$, $\nu_{\mathrm{v}}=1$, and $\nu_{\mathrm{d}}=10$. We find that the rotational polar solution emerges from the isotropic state in a subcritical bifurcation when varying $\nu_{\mathrm{p}}$, see Fig. 3(a). There, we plot the magnitude $v_{0}$ of the velocity as a function of $\nu_{\mathrm{p}}$.
In fact, it turns out that the two solutions in Eqs. (35)–(38) and (42), distinguished by the sign of ±$\pm A$, correspond to the stable and unstable analytical solutions associated with this subcritical bifurcation. In Fig. 3(a), stable and unstable branches are marked by solid and dashed lines, respectively. The presence of three branches of solution, one corresponding to the isotropic solution, one to the stable rotating polar solution, and one to the unstable rotating polar solution, indicates hysteretic behavior.
@@
We further investigate the subcritical transition to this rotational state by numerical calculations. To this end, we solve Eqs. (22)–(24) for small spatially periodic systems, see Appendix C for details. The results are displayed in Fig. 3(a) as the red and blue data points. In the first set of calculations, we start from an isotropic, macroscopically quiescent state. Increasing the active forcing $\nu_{\mathrm{p}}$ (red dots), the system at a certain threshold jumps to the rotational solution of finite amplitudes $P_{0}$, $v_{0}$, and $u_{0}$. However, when starting from an already rotating state and decreasing $\nu_{\mathrm{p}}$ (blue dots), the rotations persist down to a lower value of $\nu_{\mathrm{p}}$. The rotational solution vanishes at finite but smaller amplitudes $P_{0}$, $v_{0}$, and $u_{0}$. This confirms the analytically identified region of hysteresis in a subcritical bifurcation scenario. Either a macroscopically quiescent, isotropic state or a solution of continuously rotating uniform spatial translation develops in this region, depending on the initial conditions.
Next, we generalize our results from purely solid-like to viscoelastic behavior that allows for terminal flow. That is, finite values of the relaxation time $\tau_{\mathrm{d}}$ are considered. For the spatially uniform rotational solution, we find that the rotation frequency $\omega_{0}$ of the polar orientational order parameter, the velocity, and the elastic memory field depend on the strength $\nu_{\mathrm{p}}$ of active forcing and the relaxation time $\tau_{\mathrm{d}}$ of the memorized displacements. Examples for analytical solutions according to Eqs. (35)–(38) are depicted in Fig. 3(b). We recall that decreasing $\tau_{\mathrm{d}}$ moves the system from rather elastic solid-like towards more fluid-like behavior.
@@
Comparing the degree of polar orientational order for the stationary and for the rotational solution shown in Fig. 3(c), we observe very different dependencies on the relaxation time $\tau_{\mathrm{d}}$. In the stationary state, decreasing $\tau_{\mathrm{d}}$, and thus reducing elastic effects, generally leads to an increase in polar order. Contrarily, the rotating state exhibits a higher degree of polar order with increasing $\tau_{\mathrm{d}}$. However, for larger activities in the rotating state, this effect diminishes and the curves in Fig. 3(c) approach each other.
Physically, sufficiently long relaxation times $\tau_{\mathrm{d}}$, that is, more pronounced elastic effects hinder the flow $\mathbf{v}$. The elastic restoring forces are too large in this case to allow for persistent migration in the same direction. As a compromise, rotations occur. Reducing the role of elasticity through decreasing $\tau_{\mathrm{d}}$ facilitates unbounded flows $\mathbf{v}$. Then, instead of rotations, persistent stationary flows combined with polar orientational order are observed.
Figure 4: State diagram for small, idealized systems that combine activity and viscoelasticity. We here impose spatial uniformity of the solutions to enable analytical considerations for illustration of the theory. Dropping this idealization leads to a reconsideration of the state diagram and novel dynamic states as described in a companion paper [47]. We hatch the areas of significant deviation when the idealization of spatial uniformity is dropped. Such deviations mainly occur when both activity and (visco)elastic effects become substantial, which corresponds to the major focus of our considerations. The diagram is displayed as a function of the relaxation time of elastic memory $\tau_{\mathrm{d}}$, which measures the importance of elastic effects, and the strength of active forcing $\nu_{\mathrm{p}}$. Green and yellow areas denote the regions of uniform stationary polar and isotropic solutions, respectively. The spatially homogeneous rotational solution exists above the red curve. A hysteretic region is indicated by the dashed black line. We set the remaining parameters to $\gamma_{\mathrm{a}}=1$, $\kappa=0$, $\eta=1$, $\mu=1$, $\nu_{\mathrm{v}}=1$, and $\nu_{\mathrm{d}}=10$.
The boundaries between the three simplified uniform solutions are further visualized in Fig. 4, which shows a state diagram in the plane spanned by the elastic relaxation time $\tau_{\mathrm{d}}$ and the strength of active forcing $\nu_{\mathrm{p}}$. To summarize, the stationary polar solution is encountered for fluid-like systems of smaller $\tau_{\mathrm{d}}$ at elevated activity, while the disordered isotropic solution is found at lower activities. The rotational solution, see Eqs. (35)–(42), emerges for more solid-like systems, that is, at larger $\tau_{\mathrm{d}}$, and at higher activity. Between the rotational solution and the two stationary states, as discussed above and in Fig. 3, a hysteretical regime is observed at the lower boundary of this area. Our analytical results and the state diagram shown in Fig. 4 refer to idealized systems of small size under periodic boundary conditions, where we enforce the solutions to remain spatially uniform. The more general case is discussed in a companion article [47], where we demonstrate that complex spatiotemporal patterns emerge in spatially extended systems. Deviations occur specifically in the regime of more pronounced activity and (visco)elastic effects, which forms our main subject in this work. We mark these areas of complex novel dynamic states by hatches in Fig. 4.
@@ VIII Conclusions
In the more fluid-like case, we find a spatially uniform state of global polar orientational order, similar to what is observed in conventional theories for nonelastic systems [17, 18, 19]. In contrast to that, solid-like properties prevent persistent global motion in one direction due to the associated elastic restoring forces. There, we observe a spatially uniform state of rotating global polar orientational order. It represents globally uniform motion in one direction, while this direction rotates. When increasing the activity starting from the disordered, isotropic state, these rotations emerge in a subcritical transition exhibiting hysteresis. Our unified description reveals that this hysteretic regime vanishes as the bifurcation changes from sub- to supercritical when decreasing the relaxation time of the memory of elastic displacements $\tau_{\mathrm{d}}$, that is, turning to more fluid-like behavior in the viscoelastic regime. Remarkably, with increasing activity, the dynamic state can transition from isotropic to stationary polar and then to rotating polar order.
Our description bridges the gap between active fluids and active solids, which have by now mostly been considered in separate frameworks. As pointed out in Ref. 74, there is a critical need for this kind of integrative theories particularly for biological media, as they may change their properties dynamically and continuously. For example, during biofilm formation, bacteria excrete extracellular polymers [5, 46, 45]. Accordingly, the film starts from still fluid-like, but over time changes its rheological properties from viscous and viscoelastic to increasingly elastic. In a settled biofilm, the polymer network forms an elastic matrix and the resulting system may then be described as an active solid [68]. On a continuum-theoretical level, our description covers such transitions between fluid- and solid-like behavior. It shall be applied in future studies accordingly.
@@
In view of possible additional extensions of our investigations, we mention the alignment of the active units with the flow field. Here, we considered the situation in which the self-propelled objects tend to align along the direction of surrounding flow $\mathbf{v}$, ($\gamma_{\mathrm{a}}>0$), because it appears most natural. The opposite case ($\gamma_{\mathrm{a}}<0$) benefits antialignment. Possibly, this situation may enhance further dynamical states, also in view of flocking and clustering transitions and active transport [75, 76, 77, 78] when we leave our assumption of low to moderate densities.
We remark that, naturally, the rotating state reminds a bit of the dynamics of chiral, self-rotating active particles [79]. However, at second glance, the underlying framework is very different. Chiral active particles by themselves tend to persistently bend their trajectories and rotate their body while migrating. Often, the necessary underlying symmetry breaking is permanent, leading to a predefinded sense of rotation. In contrast to that, rotation in our case results from spontaneous symmetry breaking. Particularly, it is an immediate consequence of the coupling between activity and (visco)elasticity. Our self-propelled objects, without perturbations, would migrate along straight trajectories. Circling is induced by the elastic restoring forces. They hinder persistent motion in one direction. As a compromise, localized circling around a fixed spot occurs. Still, chiral active particles show various remarkable phenomena [79], for example, the formation of disordered hyperuniform structures [80]. Future research may additionally investigate the behavior of chiral active units embedded in viscoelastic background media using our approach.
Further extensions concern higher densities. Then, mutual steric and other short-ranged interactions become more significant. Corresponding situations require suitable statistical considerations for developing the theoretical description by coarse-graining from the discrete particle level. Examples are dynamical density functional theories for active objects [81, 82], Boltzmann-equation-type constructions [83, 84] and related approaches [85]. Their coupling to the generalized hydrodynamic theory of viscoelasticity [62, 67, 73] is an obvious task for the future.
@@ Acknowledgements.
Data Availability Statement
The code used to generate the data in support of the findings of this article is openly available [86].
Appendix A Elastic stress and memory field of elastic displacements
To ensure clarity for all readers, particularly those who mainly work with descriptions of the Lagrange type and have not been in contact with the Euler perspective for a while, we recall some of the basic items of the well-established background [61, 62] of Eq. (20). Generally, Lagrange descriptions are trajectory-based. We fix our attention to the material elements and observe how they move. The trajectory of a material element, $\mathbf{R}(t)$, is obtained from the Euler frame by integrating over time the velocity field at its position at each moment in time, $\mathbf{R}(t)=\int_{t_{0}}^{t}\mathbf{v}\left(\mathbf{R}(t^{\prime}),t^{\prime}\right)\mathrm{d}t^{\prime}$. $t_{0}$ marks the time of the beginning of our consideration. Thus, we assign to any material element located at time $t_{0}$ at position $\mathbf{R}(t_{0})$ where we will find it later, namely, $\mathbf{R}(t)$. The distance between the two positions can be considered as displacement.
In the Euler framework, the perspective is different. We now describe the system in its the current, possibly deformed state at time $t$, with the frame of reference fixed to the physical space (not the material). Thus, for any position in space $\mathbf{x}$ that we associate with the deformed state of an elastic material, we describe from which position $\mathbf{x}-\mathbf{u}(\mathbf{x},t)$ the material element currently located at $\mathbf{x}$ at time $t$ has come from. Therefore, the field $\mathbf{u}$ can be considered as an elastic memory of the previous positions. The material elements tend to displace back to these positions, if external or internal forces causing the deformation were released.
If the system is not perfectly elastic, induced elastic stresses or strains resulting from deformations may relax over time, although the overall shape of the system is maintained in the deformed state. This may happen through internal restructuring or structural rearrangement. An illustrative example are entangled polymer melts that can relax their internal elastic stresses and strains through disentanglement processes [87] (we recall that, particularly for linear elasticity and in generalized hydrodynamics [62], elastic stresses and strains in the Euler framework are equivalent due to elastic stress-strain relations). This relaxation process is one manifestation of viscoelasticity. It means that the positions to which the material elements tend to displace back to have approached the current space positions. The former are given by $\mathbf{x}-\mathbf{u}(\mathbf{x},t)$, the latter by $\mathbf{x}$. Thus, when the former approach the latter, $\mathbf{u}$ decays over time. The decay rate is called $\tau_{\mathrm{d}}^{-1}$ in Eq. (20). We have described these relations in an illustrative way here. Mathematically, they follow in a straightforward way from the systematic generalized hydrodynamics approach [62, 67, 63, 73].
For more convenient reference, we summarize the background of Eq. (20). The corresponding systematic generalized hydrodynamics theory was derived and presented about a quarter of a century ago [62, 67]. We start with the static part of the generalized hydrodynamics theory, where we here, for brevity, confine ourselves to relations associated with (visco)elasticity. This concerns the elastic stress-strain relation. It follows from the elastic energy density $\epsilon$ in the Eulerian frame of reference, which reads
$\epsilon=\frac{1}{2}K_{ijkm}{U}^{\mathrm{nonlin}}_{ij}{U}^{\mathrm{nonlin}}_{km}\,.$ (43)
Here, the components of the nonlinear strain tensor in terms of the memory field $\mathbf{u}$ in the Eulerian frame [66, 62] are given by
${U}^{\mathrm{nonlin}}_{ij}=\frac{1}{2}\big{[}\partial_{x_{i}}u_{j}+\partial_{x_{j}}u_{i}-(\partial_{x_{i}}u_{k})(\partial_{x_{j}}u_{k})\big{]}\,.$ (44)
We remark that the minus sign becomes a plus sign in the Lagrange frame [66]. The fourth-rank tensor $\mathbf{K}$ is the stiffness tensor, which for an isotropic material is of the form